OpenStudy (anonymous):

Please help! Find the four geometric means between 4 and 9724

OpenStudy (tkhunny):

\(4\;\;4r\;\;4r^{2}\;\;4r^{3}\;\;4r^{4}\;\;9724 = 4r^{5}\)

OpenStudy (anonymous):

Can you explain please?

OpenStudy (tkhunny):

I just did. What's a "Geometric Mean"?

OpenStudy (anonymous):

The geometric mean is nth root of the product of the numbers.

OpenStudy (tkhunny):

That's pretty good, but that helps you with only ONE such mean. What does it mean when there are four?

OpenStudy (anonymous):

I don't know

OpenStudy (tkhunny):

Well, that's what I wrote. Check it out for one geometric mean between to given values. 7 _____ 23 We have established the answer as \(\sqrt{7\cdot 23} = 12.68857754\) The expression I wrote is, for this smaller problem, equivalent to \(7\;\;7r\;\;23 = 7r^{2}\) The last expression leads to \(r = 1.81265393435\) And we see: \(7\) \(7\cdot 1.81265393435 = 12.68857754\) \(7\cdot (1.81265393435)^{2} = 23\) Do you see that we managed the same result? Multiple Geometric means are numbers in Geometric Sequence. This means they MUST differ by a common ratio. We're just finding the common ratio. Using the cute square root formula for calculating one creates the geometric mean directly, without calculating the common ratio. The stuff I wrote merely calculates the common ratio and then we use it to calculate each successive Geometric Mean.

OpenStudy (tkhunny):

What did you get?

OpenStudy (anonymous):

12,36,108,324?

OpenStudy (tkhunny):

You have a common ratio of 3. 9724 / 324 = 30.012345679 -- Something Fishy, there. 324 * 3 = 972 -- Were we shooting for 972 or 9724?

OpenStudy (anonymous):

I re read the problem and my teacher forgot to put a period. "Find the four geometric means between 4 and 9724,___,___,___,___,972Enter your answers separated by a comma." So I was supposed to go from 4-972

OpenStudy (tkhunny):

ahahaha. That makes SO MUCH more sense. Good work!

OpenStudy (anonymous):

Seriously haha, thank you so much!